OpenStudy (anonymous):

How many gallons of a 80% antifreeze solution must be mixed 70 gallons of 15% antifreeze to get a mixture that is 70% anitfreeze?

OpenStudy (anonymous):

LET number OF gallons OF 80% ANTIFREEZE be X THEN 0 .15 * 70 + 0.80x = 0.70(x + 70) 0.8x - 0.70x = 49 - 10.5 0.1X = 38.5 x = 385 GALLONS

OpenStudy (anonymous):

OK guy?

OpenStudy (anonymous):

good looking out jimmy

OpenStudy (anonymous):

thanx